The style of seismicity which has been occurring over the past 11 days continued throughout the period under review. Hybrid events occurred at a frequency of three to four per minute and amplitudes were variable. The total seismic energy release peaked during the night due to the incoming of moderate broadband tremor between 18:30 on 28 January and 06:30 on 29 January which was just recorded on the Gages seismometer. This may have been due to increased steam venting in the crater area. One volcano-tectonic and one long-period event occurred during this period. Both occurred beneath the crater area; the v-t occurred at a depth of 2.8 km and the l-p at 0.5 km. The v-t had an estimated magnitude of 2 to 2.5 and was reportedly felt in the Long Ground area. A temporary three-component seismometer station was today installed at Hermitage Estate above Long Ground to enable better constraining of the location and depth of the current hybrid earthquake swarm.
EDM measurements were carried out today on the Long Ground - White's Yard - Castle Peak triangle. The results from these measurements have not yet been processed but initial indications suggest that there have been no significant changes. The reflectors on Chances Peak and Gages Wall were both cleaned and re-aligned today as readings had not been possible in the past few days.
Visual observations were attempted today from Chances Peak. Unfortunately, scientists were at the crater rim during the worst part of the day, and visibility was generally poor. Rock falls continue from the north and south ends of the lava dome, which both appear still to be growing. A helicopter flight during better visibility in the afternoon enabled scientists to observe the dome both close up and from high above it. These observations indicate that current dome growth rate is at a somewhat increased level compared with the rate over the past few weeks, a situation consistent with the current high level of seismicity. The topmost part of the dome has been visible from the Richmond Hill and Fox's Bay areas for several days and may become visible from other parts of Plymouth over the next few weeks; overall dome growth rate continues to be slow and the appearance of new material visible over the edge of the crater is no cause for concern.
Gas and acid production remains high from the crater area judging from the levels encountered at the crater rim today. However, dilution by the wind ensures that gas concentrations are well below levels of concern in any inhabited areas on the island.